# Quantum Aspects of Natural Top Quark Condensation

### v2

@@ Abstract
In top quark condensation the Brout-Englert-Higgs (BEH) boson is a $\overline{t}t$ bound state. With a UV completion of a single coloron exchange interaction, a recent semiclassical treatment, main ; main2 , gave a novel theory of the BEH boson as an extended object with composite scale $M_{0}\sim 6$ TeV. Presently we obtain the semiclassical theory as an effective action, using the source/Legendre transformation techniques of Jackiw, et al. ,\!\!, Jackiw ; CJT , and fermion loop effects in the large-$N_{c}$ limit by deploying an auxiliary field to implement the sum of leading fermion loop diagrams. The theory remains natural at the loop level, with fine tuning at the level of a few %, and the effective coupling of the 4-fermion interaction, $\overline{g}_{0}^{2}$, is significantly enhanced by quantum loops over the fundamental coloron coupling, $g_{0}^{2}$. Hence a relatively weaker “topcolor” theory can produce critical coupling in the effective BEH bound state theory.
preprint: Delafield-0727-2025
August 5, 2025
@@ I Introduction
Recently we revisited the idea of “top quark condensation,” i.e., that the Brout-Englert-Higgs (BEH) boson is composed of top + anti-top quarks main ; main2 . Our new approach describes a bound state, consisting of a pair of relativisitic chiral fermions, arising from a single particle exchange of a perturbative massive gluon-like object with mass $M_{0}$ and coupling $g_{0}$ (dubbed a “coloron”). A semiclassical analysis, involving “bilocal fields,” leads to a low-mass composite BEH boson with an extended internal wave-function, $\phi(r)$ main ; main2 . $\phi(r)$ satisfies a Schrödinger-Klein-Gordon (SKG) equation in a Yukawa potential with coupling $g_{0}^{2}$. Close to its critical coupling, $g_{0}^{2}\sim g_{c}^{2}\approx 8\pi^{2}/N_{color}$ the internal wave-function spreads, thus diluting $\phi(0)$. The BEH boson is then an extended object with a small eigenvalue, $\mu^{2}$. We identify $\mu^{2}=-(88)^{2}$ GeV2, the BEH mass in the symmetric phase of the standard model (SM). The composite scale, the mass of the $SU(3)$ octet colorons, is then found to be $M_{0}\sim 6$ TeV. Near critical coupling the low energy theory is approximatly scale invariant, in analogy to a second order phase transition in condensed matter physics. This effective scale invariance leads to the spreading of the wave-function $\phi(r)$, and is a universal phenomenon in any bound state system near its critical coupling, where the eigenvalue of the Hamiltonian is small compared to the internal parameters of the theory. Hence scale symmetry is realized dynamically as the custodial symmetry of the small mass BEH bound state.
@@ II.1 Sources and Legendre Transformation
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{TC}=\!\int\!d^{4}x\biggl{(}[\overline{\omega}^{i}_{L}(x)\;iD\!\!\!/\!\,\,\omega_{iL}(x)]+[\overline{\omega}_{R}(x)\;iD\!\!\!/\!\,\,\omega_{R}(x)]+g_{0}^{2}\!\int\!d^{4}x\;d^{4}y\;[\overline{\omega}^{i}_{L}(x)\omega_{R}(y)]D_{F}(x-y)[\overline{\omega}_{R}(y)\omega_{iL}(x)]$ (5)
@@
$\!\!\!\!\!\!\!\!\!\!S_{TC}=\!\int\!d^{4}x\biggl{(}[\overline{\omega}^{i}_{L}(x)\;iD\!\!\!/\!\,\,\omega_{iL}(x)]+[\overline{\omega}_{R}(x)\;iD\!\!\!/\!\,\,\omega_{R}(x)]\biggr{)}$
@@
$S_{TC}=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)},$ (9)
@@
$\!\!\!\!\!\!\!\!\!\!S_{TC}(\Xi,\Omega)=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}+\!\int\!d^{4}x\biggl{(}\overline{\Xi}(x)\Omega(x)+\overline{\Omega}(x)\Xi(x)\biggr{)},$
$\qquad\qquad\qquad\qquad\qquad\qquad\makebox{where:}\qquad\qquad{\Xi}\equiv\begin{pmatrix}\xi_{iR}\\ \xi_{L}\end{pmatrix},\qquad\overline{\Xi}\equiv\begin{pmatrix}\overline{\xi}^{i}_{R}&\overline{\xi}_{L}\end{pmatrix}$. (11)
@@
$W=-i\hbar\ln\!\!\int\!\!D\Omega D\overline{\Omega}\;\exp\biggl{(}\frac{i}{\hbar}S^{\prime}(\Xi,\Omega)\biggr{)}$ (12)
@@
$S_{TC}=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-\overline{\Xi}(x)K^{-1}(x,y)\Xi(y)-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}$ (14)
@@
We then define the effective semiclassical field, $\Psi$:
@@
$\overline{\Psi}(x)=\frac{\delta W(\Xi)}{\delta\Xi(x)}=-\!\int\!\!d^{4}y\;\overline{\Xi}(y)K^{-1}(y,x);\qquad\Psi(x)=\frac{\delta W(\Xi)}{\delta\overline{\Xi}(x)}=-\!\int\!\!d^{4}y\;K^{-1}(x,y)\Xi(y);$ (17)
@@
We then perform a Legendre transformation to obtain the effective action, $S(\Psi,B)$, swapping $\Xi$ for the semiclassical field, $\psi$, retaining the auxiliary field $B$:
@@
$\qquad\qquad+\!\int\!d^{4}x\;d^{4}y\;\left(g_{0}D_{F}(x-y)(B^{i\dagger}(x,y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]+[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]B_{i}(x,y)\right)$ (20)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{TC}=\!\int\!d^{4}x\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]+g_{0}^{2}\!\int\!d^{4}x\;d^{4}y\;[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)].$ (22)
@@ II.2 Full Effective Action for Bilocal BEH boson via Cornwall-Jackiw-Tomboulis CJT
$S=W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}$
$+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+\left(g_{0}(B^{i\dagger}+Q^{i\dagger})D_{F}[\overline{\psi}_{R}\psi_{iL}]_{b}+g_{0}B^{i\dagger}D_{F}[\overline{\psi}_{R}\psi_{iL}]_{f}+h.c.\right)\biggr{)}$ (23)
Here we abbreviate integrals as:
$\int_{u...v}\!\!\!\!=\int\!\!d^{4}u..d^{4}v\;$ (24)
(and we will suppress writing field arguments where obvious, e.g., $B^{i}{}^{\dagger}(x,y)D_{F}(x,y)B_{i}(x,y)\rightarrow B^{i}{}^{\dagger}D_{F}B_{i}$, etc.).
@@
$\frac{\delta S}{\delta Q^{i\dagger}}=g_{0}D_{F}[\overline{\psi}_{R}\psi_{iL}]_{b}\equiv g_{0}\sqrt{N_{c}}M_{0}^{2}D_{F}H_{i};\qquad\frac{\delta S}{\delta Q_{i}}=g_{0}D_{F}[\overline{\psi}^{i}_{L}\psi_{R}]_{b}\equiv g_{0}\sqrt{N_{c}}M_{0}^{2}D_{F}H^{i\dagger};$ (25)
@@
We perform the Legendre transformation, following CJT , and using eq.(25):
$S^{\prime}=S-\int_{xy}\biggl{(}Q^{\dagger}\frac{\delta S}{\delta Q^{\dagger}}+Q\frac{\delta S}{\delta Q}\biggr{)}$ (27)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S^{\prime}=W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+g_{0}(B^{i\dagger}D_{F}C_{i}+h.c.)\biggr{)}$ (28)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!B_{i}(x,y)=g_{0}C_{i}(x,y),\qquad B^{i\dagger}(x,y)=g_{0}C^{i\dagger}(x,y)$ (30)
@@
$S^{\prime}(\Psi,H)=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]\biggr{)}$
@@
$\qquad\qquad+[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}\biggr{)}$ (31)
@@ II.3 Kinetic Terms and Factorization Ansatz
$S_{KT}=\!\int_{x}\!\biggl{(}[\overline{\psi}^{i}_{L}(x)iD\!\!\!/\!\,\,\psi_{iL}(x)]_{b}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{b}\biggr{)}$ (32)
These have to be replaced by the kinetic term of the bound state field $H_{i}(x,y)$. From eq.(25), in the free field limit (small $g_{0}$), the kinetic terms $S_{KT}$ imply the equations of motion. Technically with $H\equiv H_{i}\frac{\tau^{i}}{2}$ these are best written as commutators,
$[D_{y},[D_{y},H(x,y)]]=0\qquad[D^{\dagger}_{x},[D^{\dagger}_{x},H(x,y)]]=0$ (33)
@@
$\!\!\!\!\!\!\!\!\!\!S_{KT}\rightarrow M_{0}^{4}\int\!d^{4}x\;d^{4}y\;\biggl{(}Z|D_{R}^{\dagger}H(x,y)|^{2}+Z|D_{L}H(x,y)|^{2}\biggr{)}$ (34)
@@
To work in the rest frame as in main we follow Yukawa, Yukawa , and for the ground state we consider a factorized ansatz for $H_{i}(X,r)$:
@@
$\!\!\int\!\!d^{4}Xd^{4}r\;ZM_{0}^{4}\biggl{(}|D_{H}H(X)|^{2}|\phi(r)|^{2}+|H(X)|^{2}|\partial_{r^{\mu}}\phi(r)|^{2}\biggr{)},$ (39)

### v3

@@ Abstract
In top quark condensation the Brout-Englert-Higgs (BEH) boson is a $\overline{t}t$ bound state. With a UV completion of a single coloron exchange interaction a recent semiclassical treatment, main ; main2 , gave a novel theory of the BEH boson as an extended object with composite scale $M_{0}\sim 6$ TeV. Presently we obtain the semiclassical theory as an effective action using the source/Legendre transformation techniques of Jackiw, et al. ,\!\!, Jackiw ; CJT , and fermion loop effects in the large-$N_{c}$ limit by deploying an auxiliary field to implement the sum of leading fermion loop diagrams. The theory remains natural at the loop level with fine tuning at the level of a few %, and the effective coupling of the 4-fermion interaction, $\overline{g}_{0}^{2}$, is significantly enhanced by quantum loops over the fundamental coloron coupling, $g_{0}^{2}$. Hence a relatively weaker “topcolor” theory can produce critical coupling in the effective BEH bound state theory.
preprint: Delafield-0820-2025
August 21, 2025
@@ I Introduction
Recently we revisited the idea of “top quark condensation,” i.e., that the Brout-Englert-Higgs (BEH) boson is composed of top + anti-top quarks main ; main2 . Our new approach describes a bound state, consisting of a pair of relativisitic chiral fermions, arising from a single particle exchange of a perturbative massive gluon-like object with mass $M_{0}$ and coupling $g_{0}$ (dubbed a “coloron”). A semiclassical analysis, involving “bilocal fields,” leads to a low-mass composite BEH boson with an extended internal wave-function, $\phi(r)$ main ; main2 . $\phi(r)$ satisfies a Schrödinger-Klein-Gordon (SKG) equation in a Yukawa potential with coupling $g_{0}^{2}$. Close to its critical coupling, $g_{0}^{2}\sim g_{c}^{2}\approx 8\pi^{2}/N_{color}$ the internal wave-function spreads, thus diluting $\phi(0)$. The BEH boson is then an extended object with a small eigenvalue, $\mu^{2}$. We identify $\mu^{2}=-(88)^{2}$ GeV2, the BEH mass in the symmetric phase of the standard model (SM). The composite scale, the mass of the $SU(3)$ octet colorons, is then found to be $M_{0}\sim 6$ TeV. Near critical coupling the low energy theory is approximately scale invariant, in analogy to a second order phase transition in condensed matter physics. This effective scale invariance leads to the spreading of the wave-function $\phi(r)$, and is a universal phenomenon in any bound state system near its critical coupling, where the eigenvalue of the Hamiltonian is small compared to the internal parameters of the theory. Hence scale symmetry is realized dynamically as the custodial symmetry of the small mass BEH bound state.
@@ II.1 Sources and Legendre Transformation
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S(\omega)=\!\int\!d^{4}x\biggl{(}[\overline{\omega}^{i}_{L}(x)\;iD\!\!\!/\!\,\,\omega_{iL}(x)]+[\overline{\omega}_{R}(x)\;iD\!\!\!/\!\,\,\omega_{R}(x)]+g_{0}^{2}\!\int\!d^{4}x\;d^{4}y\;[\overline{\omega}^{i}_{L}(x)\omega_{R}(y)]D_{F}(x-y)[\overline{\omega}_{R}(y)\omega_{iL}(x)],$ (5)
@@
$\!\!\!\!\!\!\!\!\!\!S(\omega,B)=\!\int\!d^{4}x\biggl{(}[\overline{\omega}^{i}_{L}(x)\;iD\!\!\!/\!\,\,\omega_{iL}(x)]+[\overline{\omega}_{R}(x)\;iD\!\!\!/\!\,\,\omega_{R}(x)]\biggr{)}$
@@
$S(\Omega,B)=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)},$ (9)
@@
$\!\!\!\!\!\!\!\!\!\!S(\Xi,\Omega,B)=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}+\!\int\!d^{4}x\biggl{(}\overline{\Xi}(x)\Omega(x)+\overline{\Omega}(x)\Xi(x)\biggr{)},$
$\qquad\qquad\qquad\qquad\qquad\qquad\makebox{where:}\qquad\qquad{\Xi}\equiv\begin{pmatrix}\xi_{iR}\\ \xi_{L}\end{pmatrix},\qquad\overline{\Xi}\equiv\begin{pmatrix}\overline{\xi}^{i}_{R}&\overline{\xi}_{L}\end{pmatrix}$, (11)
@@
$W=-i\hbar\ln\!\!\int\!\!D\Omega D\overline{\Omega}\;\exp\biggl{(}\frac{i}{\hbar}S(\Xi,\Omega,B)\biggr{)}.$ (12)
@@
$S(\Xi,\Omega,B)=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-\overline{\Xi}(x)K^{-1}(x,y)\Xi(y)-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}.$ (14)
@@
where the quantum loops are contained in $W(\Omega)$. We then define the effective semiclassical field, $\Psi$:
@@
$\overline{\Psi}(x)=\frac{\delta W(\Xi)}{\delta\Xi(x)}=-\!\int\!\!d^{4}y\;\overline{\Xi}(y)K^{-1}(y,x);\qquad\Psi(x)=\frac{\delta W(\Xi)}{\delta\overline{\Xi}(x)}=-\!\int\!\!d^{4}y\;K^{-1}(x,y)\Xi(y).$ (17)
@@
We then perform a Legendre transformation to obtain the effective action, $S(\Psi,B)\equiv W$, swapping $\Xi$ for the semiclassical field, $\psi$, retaining the auxiliary field $B$:
@@
$\qquad\qquad+\!\int\!d^{4}x\;d^{4}y\;\left(g_{0}D_{F}(x-y)(B^{i\dagger}(x,y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]+[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]B_{i}(x,y)\right).$ (20)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S(\Psi)=\!\int\!d^{4}x\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]+g_{0}^{2}\!\int\!d^{4}x\;d^{4}y\;[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)].$ (22)
@@ II.2 Full Effective Action for Bilocal BEH boson via Cornwall-Jackiw-Tomboulis CJT
$S(\Psi,B,Q)=W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}$
$+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+\left(g_{0}(B^{i\dagger}+Q^{i\dagger})D_{F}[\overline{\psi}_{R}\psi_{iL}]_{b}+g_{0}B^{i\dagger}D_{F}[\overline{\psi}_{R}\psi_{iL}]_{f}+h.c.\right)\biggr{)},$ (23)
where,
$\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}=\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{b}+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]_{b}+[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]_{f}\biggr{)}.$ (24)
(Here we abbreviate integrals as:
$\int_{u...v}\!\!\!\!=\int\!\!d^{4}u..d^{4}v\;,$ (25)
and we will suppress writing field arguments where obvious, e.g., $B^{i}{}^{\dagger}(x,y)D_{F}(x,y)B_{i}(x,y)\rightarrow B^{i}{}^{\dagger}D_{F}B_{i}$, etc.).
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\delta S(\Psi,B,Q)}{\delta Q^{i\dagger}}=g_{0}D_{F}[\overline{\psi}_{R}\psi_{iL}]_{b}\equiv g_{0}\sqrt{N_{c}}M_{0}^{2}D_{F}H_{i};\qquad\frac{\delta S(\Psi,B,Q)}{\delta Q_{i}}=g_{0}D_{F}[\overline{\psi}^{i}_{L}\psi_{R}]_{b}\equiv g_{0}\sqrt{N_{c}}M_{0}^{2}D_{F}H^{i\dagger};$ (26)
@@
We perform the Legendre transformation, where $Q$ is now viewed as a functional of $H$ by eq.(26), following CJT :
$S(\Psi,B,H)=S(\Psi,B,Q)-\int_{xy}\biggl{(}Q^{\dagger}\frac{\delta S}{\delta Q^{\dagger}}+Q\frac{\delta S}{\delta Q}\biggr{)},$ (28)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S(\psi,B,H)=W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+g_{0}(B^{i\dagger}D_{F}C_{i}+h.c.)\biggr{)},$ (29)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!B_{i}(x,y)=g_{0}C_{i}(x,y),\qquad B^{i\dagger}(x,y)=g_{0}C^{i\dagger}(x,y).$ (31)
@@
$S(\Psi,H)=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{b}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{b}+[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr{)}$
@@
$\qquad\qquad+[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}\biggr{)}.$ (32)
@@ II.3 Kinetic Terms and Factorization Ansatz
$S_{KTb}=\!\int_{x}\!\biggl{(}[\overline{\psi}^{i}_{L}(x)iD\!\!\!/\!\,\,\psi_{iL}(x)]_{b}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{b}\biggr{)},$ (33)
These have to be replaced by the kinetic term of the bound state field $H_{i}(x,y)$. From eq.(26), in the free field limit (small $g_{0}$), the kinetic terms $S_{KTb}$ imply the equations of motion. Technically with $H\equiv H_{i}\frac{\tau^{i}}{2}$ these are best written as commutators,
$[D_{y},[D_{y},H(x,y)]]=0\qquad[D^{\dagger}_{x},[D^{\dagger}_{x},H(x,y)]]=0.$ (34)
@@
$\!\!\!\!\!\!\!\!\!\!S_{KTb}\rightarrow S_{KTH}=M_{0}^{4}\int\!d^{4}x\;d^{4}y\;\biggl{(}Z|D_{R}^{\dagger}H(x,y)|^{2}+Z|D_{L}H(x,y)|^{2}\biggr{)},$ (35)
@@
To work in the rest frame as in main we follow Yukawa, Yukawa , and for the ground state we consider a factorized ansatz for $H_{i}(X,r)$: [^4] Note that we have not specified that $H(x,y)$ or $Q(x,y)$ is factorized in $(x,y)$, as $\sim\overline{\psi}_{R}(x)\psi_{L}(y)$, hence we are permitted to make the ansatz $H_{i}(X)\phi(r)$. An $(X,r)$ factorized form is a mixture of $(x,y)$ factorized basis functions, e.g., $H(X)\phi(r)\sim\sum_{ij}a_{ij}\overline{\psi}_{Ri}(x)\psi_{Lj}(y)$, i.e., the boundstate is an entangled combination of free scattering states.
@@
$S_{KTH}\rightarrow\!\!\int\!\!d^{4}Xd^{4}r\;ZM_{0}^{4}\biggl{(}|D_{H}H(X)|^{2}|\phi(r)|^{2}+|H(X)|^{2}|\partial_{r^{\mu}}\phi(r)|^{2}\biggr{)},$ (40)